# Benchmarking Linguistic Diversity of Large Language Models

### v1

@@ 4 Settings for Diversity Benchmarking
Large language models. We evaluate the following families of models: Llama (Dubey et al., 2024), Mistral (Jiang et al., 2023), Olmo (Groeneveld et al., 2024), Gemma (Team et al., 2024), Qwen (Yang et al., 2024), Falcon (Almazrouei et al., 2023). To ensure comparability, we select the latest version of each model family that is closest in scale to 7 billion parameters. The scale selected for each model is specified in the legend of Table 1. We purposefully include models developed by organizations from different countries to be culturally inclusive. For language modeling, we use base models. For all other tasks, we employ instruction-tuned versions.
5 Results for Diversity Benchmarking
Figure 1 presents the benchmarking results of linguistic diversity across various tasks. Round dots represent the diversity of model outputs, while solid lines represent human reference outputs. Dashed lines depict the diversity of task-specific inputs (as detailed in Table 1), reflecting the conditions under which the outputs were generated. Tasks are organized in ascending order of human reference diversity for each aspect of diversity. For the machine translation task, the inputs are in French; hence, semantic diversity is measured using a multilingual SentenceBERT (Reimers and Gurevych, 2020), and syntactic diversity is evaluated with a French-specific dependency parser. As a result, these scores may not be directly comparable to those for English. We analyze the results in Figure 1 in Sections 5.2 and 5.3.
In this section, we first analyze metric correlations in Section 5.1, then compare diversity scores across tasks and models in Section 5.2. Finally, we perform a case study on syntactic diversity in story generation, comparing human and model outputs in Section 5.3.
@@ 5.1 Correlation Study
Correlation between diversity and quality. As noted in Section 2.1, diversity matters only if the text is of plausible quality. Figure 3 illustrates the correlation between diversity and quality in model outputs, using task-specific automatic metrics as quality indicators. For the language modeling task, perplexity is used to evaluate the model’s performance on reference text continuations. For machine translation, we use COMET (Rei et al., 2020), which takes into account both the source text and reference translation. For the remaining three tasks, BERTScore (Zhang et al., 2020) is used to measure the relevance between inputs and outputs. However, for these last three tasks with a subjective nature, this relevance score serves only as a proxy for quality, as automatic metrics for such tasks generally exhibit low correlation with human judgments (Chhun et al., 2024; Liu et al., 2023).
Our results show a positive correlation between quality and lexical as well as semantic diversity in model outputs. In contrast, syntactic diversity often exhibits negative correlations, where higher syntactic diversity is associated with lower quality scores. This may be attributed to the tested domains inherently exhibiting low ground-truth syntactic diversity (e.g., in language modeling) or to the limitations of quality metrics in recognizing the value of syntactic variation (e.g., in summarization, automatic story generation, and next utterance generation). These findings highlight the need to report diversity metrics alongside quality metrics for comprehensive evaluation.
@@
Llama Mistral Qwen Gemma Falcon OLMo
Precision 99.20 99.20 99.47 99.07 99.63 99.73
Recall 35.20 65.87 75.27 37.97 75.00 39.40
Table 3: Comparison of dependency tree distributions between humans and models.
@@ 5.2 Comparison Across Tasks and Models
Human output diversity. Human-level diversity varies across tasks, with no clear correlation observed among different aspects. Notably, utterances in human dialogue exhibit the lowest lexical diversity and the highest syntactic diversity, unlike the written text present in the remaining four categories. The low lexical diversity may be attributed to the conversations being specifically scripted for English learners to practice daily-life dialog. These dialogs focus on generic topics, leading to a limited range of vocabulary. In contrast, the high syntactic diversity can be explained by the inherent spontaneity of conversational language, where different speakers tend to vary significantly in their use of syntactic structures (Healey et al., 2014; Dubuisson Duplessis et al., 2017).
Model output diversity. LLMs generally lack diversity compared to humans, especially remarkable for tasks demanding high levels of creativity, such as story generation. Overall, the scores of different LLMs across tasks and diversity aspects tend to resemble each other due to the use of similar development procedures, architectures, and datasets. Compared to human outputs, LLMs generally exhibit less diversity by a considerable margin, with only a few exceptions. This is especially evident in the task of story generation, which demands the highest levels of creativity and freedom of expression, LLMs consistently lag behind humans in all three diversity aspects. In contrast, for tasks like next utterance generation, LLMs surpass human references in both lexical and semantic diversity. This discrepancy arises because the DailyDialog dataset focus on generic, everyday topics designed for English language learning, while LLMs, unconstrained by this context, frequently steer conversations toward more complex topics.
@@ 5.3 Comparing Syntactic Diversity Between Humans and Models
We adopt the Precision-Recall framework proposed by Le Bronnec et al. (2024). In our study, we substitute the original GPT-2 embeddings with the distribution of dependency trees, allowing for a fully interpretable analysis of syntactic structural differences between human and model outputs. Precision thus quantifies the percentage of model-generated dependency trees that fall within the neighborhood of human-written ones, while recall measures the percentage of human-written dependency trees within the neighborhood of model-generated ones. The method for computing the distance matrix between dependency trees is detailed in Section 3.1.1. All other hyper-parameters remain consistent with the original work (Le Bronnec et al., 2024).
Table 3 presents the precision and recall scores for all evaluated models on the story generation task. The results reveal that all models exhibit near-perfect precision, indicating that almost all generated sentences are syntactically plausible. However, the recall scores are significantly lower across all models, highlighting their inability to cover the full diversity of human syntax. This points to a notable gap between models and humans in syntactic diversity for the story generation task requiring high creativity.
To further illustrate these findings, Table 2 lists examples of syntactic patterns (POS tag n-grams) that are frequently found in human dependency trees but are missing from the model-generated ones. Conversely, we also identify syntactic patterns that models over-generate but are less common in human outputs. Recent studies (Shaib et al., 2024b) indicate that models often memorize syntactic templates encountered during pretraining, which are rarely overwritten during fine-tuning. This suggests that the observed gap in syntactic patterns may stem from a mismatch between pretraining and downstream task domains.
@@ 6 Factors Influencing LLM Diversity
In this section, we explore key factors that may influence the diversity of LLM outputs. The factors under consideration include decoding parameters, pretraining token counts, instruction tuning, model scale, and quantization. For decoding parameters and instruction tuning, we conduct experiments across all models. We employ OLMo for assessing the impact of pretraining token counts, which provides full access to its pretraining datasets, training code, and model weights at various checkpoints throughout its development. Since OLMo models are available in only two sizes, we additionally leverage Qwen2.5 models (Yang et al., 2024) to investigate the effects of model scale and quantization.
All experiments in this section are conducted for story generation, selected for its minimal constraints and high emphasis on creativity, making it an ideal benchmark for linguistic diversity. Additionally, as illustrated in Figure 1, we observe that all models substantially underperform relative to human scores in terms of diversity metrics on the story generation task.
@@ 6.1 Impact of Training Stages
OLMo was pretrained on the Dolma corpus (Soldaini et al., 2024) before going through supervised fine-tuning (SFT) on Tulu v2 (Ivison et al., 2023) and direct preference optimization (DPO) (Rafailov et al., 2023) on Ultrafeedback (Cui et al., 2024). DPO is performed on top of the model that underwent SFT. We take different checkpoints throughout the pretraining stage as well as the models after SFT and DPO. We use them to perform the story generation task and analyse the linguistic diversity of outputs. The results are presented in Figure 4. Initially, lexical diversity is exceptionally high, as expected for an untrained model that generates random tokens. This metric drops sharply after the first checkpoint (2B tokens) but then gradually increases throughout the pretraining process, without reaching saturation. In contrast, syntactic diversity also experiences a sharp decline early on; however, it saturates much more quickly, fluctuating within a narrow range afterward. Semantic diversity shows a steady increase from the beginning but also saturates relatively quickly. These observations suggest that while increasing training data generally improves lexical diversity, alternative strategies are needed to enhance syntactic and semantic diversity.
In the later stages, beyond pretraining, SFT has minimal impact on any diversity metric, while DPO leads to a decrease in syntactic diversity and an increase in lexical diversity. We now explore the impact of instruction tuning across all models in greater detail.
Impact of instruction tuning. To complement the previous discussion, we compare the output diversity between the base versions and instruction-tuned versions of all models in the context of story generation. The results, presented in Figure 6, reveal a consistent pattern across models: instruction-tuned versions show higher lexical diversity compared to their base counterparts but exhibit reductions in syntactic and semantic diversity. Notably, the decline in syntactic diversity is more pronounced than that in semantic diversity. These findings indicate that while additional training—regardless of the stage—enhances vocabulary richness, aligning models with human preferences tends to constrain them to a narrower range of grammatical structures and meanings.
@@ 6.2 Impact of Decoding Parameters
Achieving a balance between quality and diversity in LLM outputs is a known challenge, as there is often a trade-off between these two aspects (Caccia et al., 2020; Zhang et al., 2021). While decoding strategies significantly affect the trade-off between $n$-gram metrics and perplexity, their influence on other facets of text generation, such as semantic and syntactic diversity, remains underexplored. Here, we investigate how varying the decoding temperature affects the outputs in the story generation task, with results visualized in Figure 5. Output quality is estimated based on their relevance to the inputs, using BERTScore as a metric.
The results show that increasing the temperature—making decoding less restrictive—leads to greater lexical diversity, with only a minor reduction in relevance to the prompts. It might be due to the creative nature of the stroy generation task that the quality-diversity trade-off is so subtle. For syntactic diversity, while most models show fluctuating performance within a certain range, some exhibit a clear downward trend, specially OLMo and Falcon, which are trained on significantly fewer tokens compared to the other models. However, no consistent trends are observed for semantic diversity metric as decoding parameters change. This aligns with the observations of Tevet and Berant (2021), which indicate that adjusting “decoding parameters” tends to affect the form of the text rather than its meaning.
Figure 6: Impact of instruction tuning.
Furthermore, we note that, across most models, the relative ranking of diversity scores remains stable as the temperature varies. This suggests that conducting experiments with a fixed temperature is sufficient for consistent evaluation. Based on our findings, we set the temperature to 0.6 for all other experiments. Figure 5 shows that at a temperature of 0.6, the relevance to prompts remains relatively high while diversity scores significantly improve compared to lower temperatures. In fact, the creators of Llama-3.1 (Dubey et al., 2024) also identified a temperature of 0.6 as achieving the optimal balance between creativity and coherence.
@@ 6.3 Impact of Model Scale and Quantization
The Qwen2.5 model has been released in various sizes, ranging from 0.5B to 72B parameters. Due to computational resource constraints, we limit our exploration of linguistic diversity to models up to 32B parameters. The results are presented in Figure 7. We observe that lexical diversity consistently increases with model size, while semantic diversity remains stable throughout. In contrast, syntactic diversity remains relatively stable overall but exhibits an initial increase followed by a decline, peaking at 7B parameters, indicating that scaling up is not always the solution to higher diversity.
Figure 7: Impact of model scale and quantization.
@@ 7 Conclusion
Our study offers crucial insights into the linguistic diversity of current LLMs. By leveraging a comprehensive evaluation framework focused on lexical, syntactic, and semantic diversity, we provide a fresh perspective beyond traditional quality metrics. Our analysis reveals that, despite the impressive capabilities of LLMs in generating coherent and contextually appropriate text, there is a significant gap when it comes to replicating the linguistic richness characteristic of human language, especially for more creative tasks. Specifically, we observe that factors such as model scale, training data volume, and fine-tuning techniques critically influence diversity metrics.
These findings raise an important concern: as LLMs become more prevalent in content creation, their outputs may trend towards homogenization, risking a loss of linguistic richness. Notably, while instruction tuning improves lexical diversity, it constrains syntactic and semantic diversity, indicating a narrowing of expressive flexibility. Our research highlights the necessity of a more holistic and forward-looking approach in developing language models, one that prioritizes the preservation of linguistic diversity alongside optimizing performance metrics. We emphasize the need for novel strategies to balance the trade-offs between diversity and quality, ensuring that future models are capable of not only mimicking human language fluency but also maintaining its inherent diversity.

### v2

@@ 4.1 Generation Tasks
4.2 Language Models
We evaluate the following families of models: Llama (Dubey et al., 2024), Mistral (Jiang et al., 2023), Olmo (Groeneveld et al., 2024), Gemma (Team et al., 2024), Qwen (Yang et al., 2024) and Falcon (Almazrouei et al., 2023). The comparison of these models across various key characteristics is provided in Table 4 in Appendix A.
To ensure comparability, we select the latest version of each model family that is closest in scale to 7 billion parameters. The scale selected for each model is specified in the legend of Figure 1. We purposefully include models developed by organizations from different countries to be culturally inclusive. For language modeling, we use base models. For all other tasks, we employ instruction-tuned versions.
5 Results of Diversity Benchmarking
Figure 1 visualizes the benchmarking results of linguistic diversity across various tasks. Round dots represent the diversity of model outputs, while solid lines represent human reference outputs. Dashed lines depict the diversity of task-specific inputs (as detailed in Table 1), reflecting the conditions under which the outputs were generated. Tasks are organized in ascending order of creativity level. The detailed numerical results are provided in Table 5 in Appendix B.
For the machine translation task, the inputs are in French; hence, semantic diversity is measured using a multilingual SentenceBERT (Reimers and Gurevych, 2020), and syntactic diversity is evaluated with a French-specific dependency parser. As a result, these scores may not be directly comparable to those for English. The diversity of human reference outputs serves as a baseline for interpreting whether the model under or over represents the diversity for each task.
In this section, we first analyze metric correlations in Section 5.1, then compare diversity scores across tasks and models in Section 5.2. Finally, in Section 5.3, we perform a case study on syntactic diversity in story generation, comparing human and model outputs.
@@ 5.1 Correlation Study
Correlation between diversity and quality. We manually verify that all models produce plausible and coherent text that meets the basic requirements for diversity evaluation across all tasks. Building on this, we examine more specific qualities of the model outputs. Figure 3 illustrates the correlation between diversity and quality in model outputs, using task-specific automatic metrics as quality indicators. For the language modeling task, perplexity is used to evaluate the model’s performance on reference text continuations. For machine translation, we use COMET (Rei et al., 2020), which takes into account both the source text and reference translation. For the remaining three tasks, BERTScore (Zhang et al., 2020) is used to measure the relevance between inputs and outputs. Due to the inherently subjective nature of these tasks, automatic metrics generally exhibit weak correlations with human judgments (Liu et al., 2023) and should be interpreted cautiously. Nonetheless, we adopt BERTScore as an approximate quality indicator, as embedding-based metrics of this kind demonstrate the strongest system-level correlation with human evaluations among available automatic measures (Chhun et al., 2022).
Our results show a positive correlation between quality and lexical as well as semantic diversity in model outputs. In contrast, syntactic diversity often exhibits negative correlations, where higher syntactic diversity is associated with lower quality scores. This may be attributed to the tested domains inherently exhibiting low ground-truth syntactic diversity (e.g., in language modeling) or to the limitations of quality metrics in recognizing the value of syntactic variation (e.g., in summarization, automatic story generation, and next utterance generation). These findings highlight the need to report diversity metrics alongside quality metrics for comprehensive evaluation, as the relationship between the two is not consistent across tasks or aspects.
@@
Llama Mistral Qwen Gemma Falcon OLMo
Precision 99.20 99.20 99.47 99.07 99.63 99.73
Recall 35.20 65.87 75.27 37.97 75.00 39.40
Table 2: Comparison of dependency tree distributions between humans and models for the story generation task.
@@ 5.2 Comparison Across Tasks and Models
Human output diversity. Human-level diversity varies across tasks, with no clear correlation observed among different aspects. Notably, utterances in human dialogs exhibit the lowest lexical diversity and the highest syntactic diversity, unlike the written text present in the remaining four tasks. The low lexical diversity may be attributed to the conversations being specifically scripted for English learners to practice daily-life dialog. These dialogs focus on generic topics, leading to a limited range of vocabulary. In contrast, the high syntactic diversity can be explained by the inherent spontaneity of conversational language, where different speakers tend to vary significantly in their use of syntactic structures (Healey et al., 2014; Dubuisson Duplessis et al., 2017). Human summaries show limited lexical and syntactic variation but exhibit the highest semantic diversity, suggesting a narrow range in form but a broad range in content. In contrast, human translations score lowest in semantic diversity, reflecting the restricted topical scope of the source texts. Wikipedia-based language modeling displays high topic diversity, while human-written stories tend to be diverse across all three dimensions.
Model output diversity. LLMs lack diversity compared to humans for tasks demanding high levels of creativity, such as story generation. Overall, the scores of different LLMs across tasks and diversity aspects tend to resemble each other, potentially due to the use of similar development procedures, architectures, and datasets. However, this remains an assumption, as most LLM developers do not fully disclose their training data or protocols, even when the models themselves are open-weight. The extent to which LLMs under- or over-represent diversity compared to humans varies significantly by the task domain. For the task of story generation, which demands the highest levels of creativity and freedom of expression, LLMs consistently lag behind humans in all three diversity aspects. In contrast, for tasks like next utterance generation, LLMs surpass human references in both lexical and semantic diversity. This discrepancy arises because the DailyDialog dataset focuses on generic, everyday topics designed for English language learning, while LLMs, unconstrained by this context, frequently steer conversations toward more complex topics.
@@ 5.3 Comparing Syntactic Diversity Between Humans and Models
We adopt the Precision-Recall framework proposed by Le Bronnec et al. (2024). This framework relies on GPT-2 embeddings, followed by Principal Component Analysis (PCA) and K-means clustering, to estimate the supports of text distributions. In our study, we replace the original GPT-2 embeddings with the implicit distribution of dependency tree embeddings induced by the WL graph kernel. Precision is defined as the proportion of dependency trees from model-generated text that lie within the support of dependency trees from human-written text. A high precision indicates that the model-generated structures are more plausible and human-like, thus reflecting their quality. Recall, on the other hand, measures the proportion of dependency trees from human-written text that fall within the support of the model-generated distribution. A high recall suggests that the model captures the full diversity of human-written structures. The method for computing pairwise distances between dependency trees is described in Section 3.1.1, which serves as the basis for constructing the distance matrix. All other hyper-parameters remain consistent with the original work (Le Bronnec et al., 2024).
Table 2 presents the precision and recall scores for all evaluated models on the story generation task. The results reveal that all models exhibit near-perfect precision, indicating that almost all generated sentences are syntactically plausible. In contrast, recall scores are substantially lower than precision scores across all models, revealing their limited capacity to capture the full breadth of human syntactic diversity. This points to a notable gap between models and humans in syntactic diversity for the story generation task where high creativity is required.
To further illustrate these findings, Table 3 lists examples of syntactic patterns (POS tag n-grams) that are frequently found in human dependency trees but are missing from the model-generated ones. Conversely, we also identify syntactic patterns that models over-generate but are less common in human outputs. Recent studies (Shaib et al., 2024b) indicate that models often memorize syntactic templates encountered during pretraining, which are rarely overwritten during SFT and preference tuning. This suggests that the observed gap in syntactic patterns may stem from a mismatch between pretraining and downstream task domains. For instance, in the pretraining corpus of OLMo, over 80% of the data originates from web pages in Common Crawl, while less than 0.3% comes from Project Gutenberg books, one of the only sources potentially aligned with the narrative style required for story generation (Soldaini et al., 2024).
@@ 6 Factors Influencing LLM Diversity
In this section, we explore key factors that may influence the diversity of LLM outputs. The factors under consideration include pretraining token counts, instruction tuning, decoding parameters, prompt formulation, model scale, and quantization. For decoding parameters, prompt formulation and instruction tuning, we conduct experiments across all models. We employ OLMo for assessing the impact of pretraining token counts since it provides full access to its pretraining datasets and model weights at various checkpoints throughout its development. Since OLMo models are available in only two sizes, we additionally leverage Qwen2.5 models (Yang et al., 2024) to investigate the effects of model scale and quantization.
All experiments in this section are conducted on the story generation task, where linguistic diversity plays a central role. Its minimal input constraints and strong emphasis on creativity make it an ideal benchmark for evaluating linguistic diversity. Moreover, as shown in Figure 1, all models fall significantly short of human performance in terms of diversity on this task, highlighting the importance of identifying which factors contribute to this gap. We emphasize that the conclusions drawn in this section are specific to the story generation task and should not be generalized to broader LLM behavior without further investigation.
@@ 6.1 Impact of Training Stages
Pretraining. We choose OLMo, pretrained on the Dolma corpus (Soldaini et al., 2024), to study the evolution of linguistic diversity during pretraining. This is because OLMo is the only model in our benchmark with publicly available intermediate checkpoints during pretraining. The results are presented in Figure 4. Intitially, lexical diversity is exceptionally high, as expected for an untrained model that generates random tokens. This metric drops sharply after the first checkpoint (2B tokens) but then gradually increases throughout the pretraining process, without reaching saturation. In contrast, syntactic diversity also experiences a sharp decline early on; however, it saturates much more quickly, fluctuating within a narrow range afterward. Semantic diversity shows a steady increase from the beginning but also saturates relatively quickly. These observations suggest that while increasing training data generally improves lexical diversity, alternative strategies are needed to enhance syntactic and semantic diversity.
Instruction tuning. We now move on to study the impact of instruction tuning on linguistic diversity. After pretraining, OLMo underwent supervised fine-tuning (SFT) on Tulu v2 (Ivison et al., 2023) and direct preference optimization (DPO) (Rafailov et al., 2023) on Ultrafeedback (Cui et al., 2024), with DPO applied on top of SFT. We observe that SFT has minimal impact on any diversity metric, while DPO leads to a decrease in syntactic diversity and an increase in lexical diversity, potentially reflecting characteristics of the SFT and DPO datasets.
Since all models in our benchmark provide both base and instruction-tuned versions, we extend our analysis to assess the impact of instruction tuning across the full set. As shown in Figure 5, the results mirror those observed for OLMo: instruction-tuned versions show higher lexical diversity compared to their base counterparts but exhibit reductions in syntactic and semantic diversity. Notably, the decline in syntactic diversity is more pronounced than that in semantic diversity. These findings indicate that while additional training—regardless of the stage—enhances vocabulary richness, aligning models with human preferences tends to constrain them to a narrower range of grammatical structures and meanings.
Figure 7: Impact of prompt formulation.
@@ 6.2 Impact of Decoding Parameters
Achieving a balance between quality and diversity in LLM outputs is a known challenge, as there is often a trade-off between these two aspects (Caccia et al., 2020). The choice of decoding strategy plays a crucial role in controlling this trade-off (Zhang et al., 2021). Here, we investigate how varying the decoding temperature affects the outputs in the story generation task, with results visualized in Figure 6. Output quality is estimated based on their relevance to the inputs, using BERTScore as a metric.
The results show that increasing the temperature—making decoding less restrictive—leads to greater lexical diversity, with only a minor reduction in relevance to the inputs. It might be due to the creative nature of the story generation task that the quality-diversity trade-off is so subtle. For syntactic diversity, while most models show fluctuating performance within a certain range, some exhibit a clear downward trend, specially OLMo and Falcon, which are trained on significantly fewer tokens compared to the other models. However, no consistent trends are observed for semantic diversity metric as decoding parameters change. This aligns with the observations of Tevet and Berant (2021), which indicate that adjusting decoding parameters tends to affect the form of the text rather than its meaning.
Furthermore, we note that, across most models, the relative ranking of diversity scores remains stable as the temperature varies. This suggests that conducting experiments with a fixed temperature is sufficient for consistent evaluation. Based on our findings, we set the temperature to 0.6 for all other experiments. Figure 6 shows that at a temperature of 0.6, the relevance to the inputs remains relatively high while diversity scores significantly improve compared to lower temperatures.
Figure 8: Impact of model scale and quantization.
6.3 Impact of Prompt Formulation
Previous studies have established that LLMs exhibit considerable sensitivity to prompt formulations, particularly affecting their performance on discriminative downstream tasks (Sclar et al., 2024; Wahle et al., 2024). Here, we explore whether the linguistic diversity of stories generated by LLMs is similarly influenced by variations in the formulation of prompts. We conduct experiments across the full range of models, and the results are depicted in Figure 7. The solid lines represent results obtained using the standard prompt, consisting of the task-specific instruction “Please continue the following story” combined with sample-specific inputs from the WritingPrompts dataset. To evaluate prompt sensitivity, we modify the prompt to explicitly encourage creativity by changing the instruction to “Please continue the following story and be as creative as possible”. Results from these modified prompts are shown as dash-dot lines in Figure 7.
Our analysis indicates that altering the prompt formulation has minimal impact on the diversity of generated stories across all three evaluated aspects. This suggests that the linguistic patterns exhibited by LLMs across creative generations represent inherent model characteristics that are less sensitive to prompt variations compared to accuracy-based performance on discriminative tasks. Consequently, enhancing linguistic diversity in LLM outputs through straightforward prompt engineering alone would be challenging.
@@ 6.4 Impact of Model Scale and Quantization
We now study the impact of model scale on linguistic diversity with the Qwen2.5 model. Qwen2.5 has been released in various sizes, ranging from 0.5B to 72B parameters. Due to computational resource constraints, we limit our exploration of linguistic diversity to models up to 32B parameters. The results are presented in Figure 8. We observe that lexical diversity consistently increases with model size, while semantic diversity remains stable throughout. In contrast, syntactic diversity remains relatively stable overall but exhibits an initial increase followed by a decline, peaking at 7B parameters, indicating that scaling up is not always the solution to higher linguistic diversity.
@@ 7 Conclusion
Our study offers crucial insights into the linguistic diversity of current LLMs. By leveraging a comprehensive evaluation framework focused on lexical, syntactic, and semantic diversity, we provide a fresh perspective beyond traditional quality metrics. Our analysis reveals that, despite the impressive capabilities of LLMs in generating coherent and plausible text, there is a significant gap when it comes to replicating the linguistic richness of human language for creative tasks such as story generation. Furthermore, we find that factors like pretraining data volume, instruction tuning, decoding strategies, model scale, and quantization significantly influence diversity metrics. In particular, while instruction tuning improves lexical diversity, it constrains syntactic and semantic diversity, indicating a narrowing of expressive flexibility. These findings raise an important concern: as LLMs become more prevalent in content creation, their outputs may trend towards homogenization, risking a loss of linguistic richness. Our research highlights the necessity of a more holistic and forward-looking approach in developing language models, one that prioritizes the preservation of linguistic diversity alongside optimizing performance metrics.